An inclined plane makes an angle of with the horizontal. Find the constant force, applied parallel to the plane, required to cause a box to slide (a) up the plane with acceleration 1.2 m/s² and (b) down the incline with acceleration 1.2 m/s². Neglect friction forces.

1
Expert's answer
2021-12-03T12:45:18-0500

"\alpha =30°"

"m=15\ kg"


(a) "F-mg\sin30°=ma\to F=ma+mg\sin30°="


"=15\cdot1.2+15\cdot9.8\cdot\sin30°=91.5\ (N)"


(b) "F+mg\sin30°=ma\to F=ma-mg\sin30°="


"=15\cdot1.2-15\cdot9.8\cdot\sin30°=-55.5\ (N)"